OpenStudy (anonymous):

Find the indefinite integral. dy/(y^4+4y^2+4) Can someone help me?

OpenStudy (fibonaccichick666):

Can you factor the square?

OpenStudy (fibonaccichick666):

for the botom

OpenStudy (fibonaccichick666):

then do you have any ideas what I might suggest next?

OpenStudy (anonymous):

\(y^4+4y^2+4=(y^2+2)^2\) so we have $$\int\frac{dy}{(y^2+2)^2}=\dots$$

OpenStudy (anonymous):

are you sure the numerator is just \(dy\)?

OpenStudy (anonymous):

yeah i'm sure

OpenStudy (fibonaccichick666):

ok, so now

OpenStudy (fibonaccichick666):

even though odrin did the factoring for you....(yes i'm looking at you) Can you think of a method where you split a denominator into two fractions?

OpenStudy (anonymous):

Splitting the denominator in 2 will not be of much use here:$$\frac1{(y^2+2)^2}=\frac{A}{y^2+2}+\frac{B}{(y^2+2)^2}\implies A=0,B=1$$

OpenStudy (anonymous):

You can salvage the partial fraction decomposition approach by agreeing to factor \(y^2+2=(y-i\sqrt2)(y+i\sqrt2)\) but you'll need to do a decent amount of simplification at the end to get 'nice' results free of any obvious imaginary numbers

OpenStudy (fibonaccichick666):

...Please stop answering questions I am posing to get the asker thinking. It is very rude

OpenStudy (anonymous):

Your suggestion was incorrect.

OpenStudy (fibonaccichick666):

AGAIN, to get them thinking

OpenStudy (anonymous):

you mean split the (y^2+2)^2 into two fractions?

OpenStudy (shubhamsrg):

Or, after converting it into (y^2 + 2)^2 , you may simply substitute y = sqrt2 * tan u , to make things easier.

OpenStudy (fibonaccichick666):

sort of, we have \[\frac{1}{(y^2+2)^2}\] we want to use something called partial fractions, are you familiar with that?

OpenStudy (fibonaccichick666):

and yea, you could, use that, but why the square root of 2?

OpenStudy (fibonaccichick666):

ah nvm

OpenStudy (fibonaccichick666):

I see, for the factoring

OpenStudy (anonymous):

i think i'm new with it (partial fractions) ><

OpenStudy (anonymous):

@shubhamsrg's suggestion was the one you were looking for @FibonacciChick666 in place of partial fraction decomposition

OpenStudy (fibonaccichick666):

ok, so if you just covered that, it is probably what they want you to use.

OpenStudy (fibonaccichick666):

So, odrin's decomp was incorrect

OpenStudy (fibonaccichick666):

can you fix it?

OpenStudy (fibonaccichick666):

This may help you look at it @dangbeau https://www.math.ucdavis.edu/~kouba/CalcTwoDIRECTORY/partialfracdirectory/PartialFrac.html

OpenStudy (anonymous):

@dangbeau In general, when you have \(y^2+a^2\) without a radical in your denominator, an appropriate trigonometric substitution like \(y=a\tan t\) is your best bet. We know this from the Pythagorean identity:$$\cos^2t+\sin^2t=1\\1+\frac{\sin^2t}{\cos^2 t}=\frac1{\cos^2 t}\\1+\tan^2 t=\sec^2 t$$

OpenStudy (anonymous):

Additionally, you may find it convenient to use a hyperbolic trigonometric substitution, as$$\cosh^2t-\sinh^2t=1\\\cosh^2t=1+\sinh^2 t$$i.e. consider \(y=\sqrt2\,\sinh t\), which yields \(dy=\sqrt2\,\cosh t\ dt\):$$\int\frac1{(y^2+2)^2}dy=\int\frac{\sqrt2\,\cosh t}{(2\sinh^2 t+2)^2}dt=\int\frac{\sqrt2\,\cosh t}{4\cosh^4 t}dt=\frac1{\sqrt8}\int\operatorname{sech}^3 t\ dt$$

OpenStudy (anonymous):

If you are not familiar with hyperbolic substitutions, this problem will likely not be any easier for you to solve, but if you're savvy with them the above might be a path of less resistance.

OpenStudy (fibonaccichick666):

@dangbeau , which approach do you like?

OpenStudy (anonymous):

hmm this is mutiple choices, and all 4 answers have arctan (y/sqrt2), can u suggest the best way for me to have that answer?

OpenStudy (fibonaccichick666):

that would be @shubhamsrg 's solution

OpenStudy (fibonaccichick666):

so long as you have covered the method of that sort of substitution, I think that would be the easiest approach. If not, I'll think on that

OpenStudy (anonymous):

i'll try

OpenStudy (fibonaccichick666):

ok, so list your steps here and we can go over and check them.

OpenStudy (anonymous):

@FibonacciChick666 pointed out there is a misleading initial step in my partial fraction decomposition earlier; what I intended was:$$\frac1{(y^2+2)^2}=\frac{Ax+B}{y^2+2}+\frac{Cx+D}{(y^2+2)^2}$$still, however, this invariably yields \(A=B=C=0,D=1\), which does not contradict the earlier post.

OpenStudy (anonymous):

\[y=\sqrt{2}tanu \]\[dy=\sqrt{2}\sec ^{2}udu\]=>\[\frac{ \sqrt{2}\sec ^{2}u }{((\sqrt{2}tanu) ^{2}+2)^{2} }\]

OpenStudy (fibonaccichick666):

good, continue

OpenStudy (anonymous):

\[\frac{ \sqrt{2}\tan^{2}u }{ (2\tan ^{2}u +2)^{2} }\] i'm stucking here T_T

OpenStudy (fibonaccichick666):

why did you change the top of the fraction to a tangent?

OpenStudy (fibonaccichick666):

also, should the derivative be on the top or bottom of the fraction?

OpenStudy (anonymous):

how to change to a tangent?

OpenStudy (fibonaccichick666):

? Sorry can you rephrase, I don't understand your question

OpenStudy (anonymous):

ah sorry i misread your sentence, i thought you told me change the top to a tangent so it's \[\frac{ \sqrt{2}\sec ^{2u}du }{(2\tan^{2}u+2)^{2}}\] what should i do next?

OpenStudy (fibonaccichick666):

so try and factor out those twos if you can

OpenStudy (anonymous):

\[\frac{ \sqrt{2}(1+\tan^{2}u )du}{ 4(\tan^{2}u+1)^{2} }\]

OpenStudy (fibonaccichick666):

now what can you do?

OpenStudy (anonymous):

i try thinking but still stuck ><

OpenStudy (fibonaccichick666):

what is the same?

OpenStudy (anonymous):

how to integrate it \[\frac{ 1 }{ \tan^{2}u+1 }du\]

OpenStudy (fibonaccichick666):

what happened to the sqrt2 and 4?

OpenStudy (anonymous):

is it put outside the integral? \[\frac{ \sqrt{2} }{ 4 }\int\limits_{}^{}\frac{ 1}{ \tan^{2}u+1 }du\]

OpenStudy (fibonaccichick666):

yea, and now you can make another sub

OpenStudy (anonymous):

can you make it more clearly?